OpenStudy (anonymous):

The sum of the squares of two positive numbers is 100; their maximum product is?

OpenStudy (anonymous):

a^2 + b^2=100...max value will be like AM>= GM

OpenStudy (anonymous):

what does your last part mean?

OpenStudy (anonymous):

arithematic mean >= geometric mean AM is \[\frac{ \sum_{1}^{n} a }{ n }\] and GM is \[ \sqrt[n]{\prod_{1}^{n} a}\]

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

The maximum product is 50. Refer to the Mathematica attachment.

OpenStudy (anonymous):

wow thanks so much!

OpenStudy (anonymous):

You are welcome.